# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ 2.2 Gaussian approximation
$\rho_{n,W}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert},\qquad\text{where }Z\sim\mathsf{N}_{d}(0,\Sigma),$
and where the dependence of $S_{n,W}$ on $c$, via $\varepsilon_{n}$, is suppressed notationally. We assume throughout that $d\geq 2$ and $n>3$ (such that, e.g., $\sqrt{\log(d)}>0$).
@@
Fix $c\in(1,\sqrt{1.5})$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5), then
$\rho_{n,W}$ $\leq C\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}$
$+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2},$ (6)
where $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
@@
Consider the case of $\overline{\eta}_{n}=0$. Theorem 2.1 then shows that winsorized means can break through the polynomial growth rate barrier $d=n^{m/2-1}$ that $d$ must obey for the Gaussian approximation error to the distribution of $S_{n}$, i.e., $\rho_{n}$ in (1), to converge to zero. In particular, $\rho_{n,W}\to 0$ if only $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which allows for $d$ growing exponentially in $n$ for any $m>2$ albeit with a small exponent for $m$ close to two.
@@
Thus, the winsorized mean obeys a Gaussian approximation result over $\mathcal{H}$ under heavy tails and adversarial contamination, which is in analogy to the Gaussian approximation result over $\mathcal{R}\subseteq\mathcal{H}$ in Resende (2024) for the trimmed mean analyzed there. However, as discussed in the introduction, the dependence of the trimmed mean estimator in Resende (2024) on $m$ implies that if the $X_{i}$ have fewer moments than the chosen $m$ then there are no approximation guarantees. If, on the other hand, the $X_{i}$ have more moments than the specified $m$ then the guarantees may be suboptimal. In contrast, the implementation of our winsorized mean does not depend on $m$ — it “adapts” to it. Furthermore, we recover the rate $d=o\mathinner{\bigl(\exp(n^{1/5})\bigr)}$ as $m\to\infty$, which is currently the best available for Gaussian approximations based on $S_{n}$ with sub-exponential $X_{i}$, instead of $d=o\mathinner{\bigl(\exp(n^{1/6})\bigr)}$ for the trimmed mean analyzed in Resende (2024). In case the entries of the $X_{i}$ are heavy-tailed in the sense of possessing exactly, e.g., $m=4$ moments, $\rho_{n,W}\to 0$ if $\log(d)=o(n^{1/9})$ whereas the trimmed mean analyzed in Resende (2024) allows $\log(d)=o(n^{3/35})$ if implemented with $m=3$, $\log(d)=o(n^{4/35})$ if implemented with $m=4$, and provides no guarantees if implemented with $m>4$. Thus, unless one has reliable information on the number of moments the data possesses (knows $m\approx 4$), our method adapting to $m$ allows for larger $d$.
@@
To prove Theorem 2.1, we first show that the order statistics $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ can (essentially) be replaced by closely related population quantiles $Q_{\varepsilon_{n},j}$ and $Q_{1-\varepsilon_{n},j}$ of the $X_{1,j}$ such that one can analyze the non-random winsorization functions $\phi_{n}(x)\mathrel{\mathop{\ordinarycolon}}=\phi_{Q_{\varepsilon_{n},j},Q_{1-\varepsilon_{n},j}}(x)$ instead of $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(x)$ for all $j=1,\ldots,d$: [^5] A similar control of the order statistics was also used in Lugosi and Mendelson (2021). Resende (2024) studied the trimmed mean by relating it to the winsorized mean.
$S_{n,W,j}\approx\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n}(\tilde{X}_{i,j})-\mu_{j}\bigr]}}_{T_{n,j}}$ $=\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n}(\tilde{X}_{i,j})-\phi_{n}(X_{i,j})\bigr]}}_{I_{n,j,1}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n}(X_{i,j})-\mathbb{E}\phi_{n}(X_{i,j})\bigr]}}_{I_{n,j,2}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\mathbb{E}\phi_{n}(X_{i,j})-\mu_{j}\bigr]}}_{I_{n,j,3}}.$
@@
The term $\max_{j=1,\ldots,d}|I_{n,j,1}|$ isolates the effect of the adversarial contamination of the data and $\max_{j=1,\ldots,d}|I_{n,j,3}|$ quantifies how far the winsorized means $\mathbb{E}\phi_{n}(X_{1,j})$ are from the population means $\mu_{j}$ of interest. Next, note that $I_{n,j,2}$ is a sample average of bounded i.i.d. random variables. Thus, letting $Z_{n}\sim\mathsf{N}_{d}(0,\Sigma_{\phi_{n}})$ with $\Sigma_{\phi_{n}}$ being the covariance matrix of $I_{n,2}=(I_{n,1,2},\ldots,I_{n,d,2})^{\prime}$, one can apply, e.g., the Gaussian approximation for sums of sub-exponential random vectors from Chernozhuokov et al. (2022) to this term to show that
@@ 3 Bootstrap approximations for winsorized means
Similarly to Gaussian approximations for $S_{n}$, cf. the references in the introduction, the one for $S_{n,W}$ in Theorem 2.1 is not directly useful for statistical inference since the covariance matrix $\Sigma$ of the approximating distribution $\mathsf{N}_{d}(0,\Sigma)$ is typically unknown. Because the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 in Chernozhuokov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)) shows that for $Z_{1}\sim\mathsf{N}_{d}(0,\Sigma^{(1)})$ and $Z_{2}\sim\mathsf{N}_{d}(0,\Sigma^{(2)})$ with $\min_{j=1,\ldots,d}\Sigma^{(2)}_{j,j}>b$ for some $b>0$, it holds that
$\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z_{1}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{2}\in H\bigr)}\Bigr\rvert}\leq C\mathinner{\Bigl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\Sigma^{(1)}_{j,k}-\Sigma^{(2)}_{j,k}\bigr\rvert}\log^{2}(d)\Bigr)}^{1/2},$ (7)
@@
In the next section we construct the estimator $\tilde{\Sigma}_{n}$ based on suitably winsorized observations $\tilde{X}_{i}$, for which we establish performance guarantees even for $\bar{\eta}_{n}>0$ and when the $X_{i}$ possess only $m>2$ moments. Our estimator does not depend on unknown population quantities and its performance guarantees “adapt” to the unknown $m$. We stress that Theorem 3.2 below is modular in the sense that it remains valid for any estimator $\hat{\Sigma}_{n}$ satisfying a bound as in Theorem 3.1.
@@ 3.1 Estimating $\Sigma$
Imposing only $m>2$ moments to exist we now construct an estimator of $\Sigma$ with precision guarantees in the maximal entrywise norm. To avoid making assumptions on $\mu$, let
$Y_{i}=\frac{1}{\sqrt{2}}\mathinner{\bigl(X_{2i}-X_{2i-1}\bigr)}\qquad\text{and}\qquad\tilde{Y}_{i}=\frac{1}{\sqrt{2}}\mathinner{\bigl(\tilde{X}_{2i}-\tilde{X}_{2i-1}\bigr)},\qquad i=1,\ldots,\lfloor n/2\rfloor.$ (8)
Clearly, $Y_{1},\ldots,Y_{\lfloor n/2\rfloor}$ are i.i.d. mean zero with covariance matrix $\Sigma$. In the sequel we assume for convenience that $n$ is even. Let $N\mathrel{\mathop{\ordinarycolon}}=n/2$ and set
$\varepsilon_{n}^{\prime}=c2\overline{\eta}_{n}+c\sqrt{\frac{\log(d^{2}N)}{2N}},\qquad c\in(1,\infty).$ (9)
Writing $\hat{a}_{j}=\tilde{Y}_{\lceil\varepsilon_{n}^{\prime}N\rceil,j}^{*}$ and $\hat{b}_{j}=\tilde{Y}_{\lceil(1-\varepsilon_{n}^{\prime})N\rceil,j}^{*}$ for $j=1,\ldots,d$, define $\tilde{\Sigma}_{n}$ as the matrix with entries
$\tilde{\Sigma}_{n,j,k}=\frac{1}{N}\sum_{i=1}^{N}\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{Y}_{i,j})\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{Y}_{i,k}),\qquad 1\leq j,k\leq d.$ (10)
$\tilde{\Sigma}_{n}$ is positive semi-definite and symmetric by virtue of being a Gram matrix and obeys the following precision guarantee.
@@
Fix $c\in(1,\infty)$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}^{\prime}$ as in (9), then for a constant $C=C(b_{2},c,m)$,
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}\leq\frac{24}{n}.$ (11)
Theorem 3.1 shows that for any $m>2$ it is possible for $\max_{1\leq j,k\leq d}\mathinner{\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\rvert}$ to converge to zero in probability even when $d$ grows exponentially in $n$. Finally, we reiterate that the approximation in Theorem 3.2 below remains valid for any estimator from the burgeoning literature on large covariance matrix estimation satisfying (11): the result is not specific to $\tilde{\Sigma}_{n}$.

### v4

@@ 2.2 Gaussian approximation
$\rho_{n,W}\mathrel{\mathop{\ordinarycolon}}=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)}\Bigr\rvert},\qquad\text{where }Z\sim\mathsf{N}_{d}(0,\Sigma).$
We assume throughout that $d\geq 2$ and $n>3$ (such that, e.g., $\sqrt{\log(d)}>0$).
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}\in(0,1/2)$, with $\varepsilon_{n}$ as in (4), then, for a constant $C$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$, we have
$\rho_{n,W}$ $\leq C\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}$ (5)
$\quad+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}.$
@@
Consider the case of $\overline{\eta}_{n}=0$. Theorem 2.1 then shows that winsorized means can break through the polynomial growth rate barrier $d=n^{m/2-1}$ that $d$ must obey for the Gaussian approximation error to the distribution of $S_{n}$, i.e., $\rho_{n}$ in (1), to converge to zero. In particular, $\rho_{n,W}\to 0$ if only $\log(d)=o\mathinner{\bigl(n^{\frac{m-2}{5m-2}}\bigr)}$, which allows for $d$ growing exponentially in $n$ for any $m>2$, albeit with a small exponent for $m$ close to two. Thus, the winsorized mean obeys a Gaussian approximation result over $\mathcal{H}$ under heavy tails and adversarial contamination, which is in analogy to the Gaussian approximation result over $\mathcal{R}\subsetneq\mathcal{H}$ in Resende (2024) for the trimmed mean analyzed there. However, let us highlight the following key differences, which we already commented on in the introduction: The implementation of the trimmed mean estimator in Resende (2024) requires knowledge of the (typically) unknown $m$. Thus, if the $X_{i}$ have fewer moments than the chosen $m$, then there are no approximation guarantees. If, on the other hand, the $X_{i}$ have more moments than the specified $m$, then the guarantees may be suboptimal. In contrast, the implementation of our winsorized mean does not depend on $m$ — it “adapts” to it. Furthermore, we recover the rate $d=o\mathinner{\bigl(\exp(n^{1/5})\bigr)}$ as $m\to\infty$, which is currently the best available for Gaussian approximations based on $S_{n}$ with sub-exponential $X_{i}$, instead of $d=o\mathinner{\bigl(\exp(n^{1/6})\bigr)}$ for the trimmed mean analyzed in Resende (2024). To make this more concrete, consider the case where the entries of the $X_{i}$ are heavy-tailed in the sense of possessing exactly, e.g., $m=4$ moments. Then, $\rho_{n,W}\to 0$ if $\log(d)=o(n^{1/9})$, whereas the trimmed mean analyzed in Theorem 2 in Resende (2024) allows $\log(d)=o(n^{3/35})$ if implemented with $m=3$, $\log(d)=o(n^{4/35})$ if implemented with the true $m=4$, and provides no guarantees if implemented with $m>4$. Thus, unless one has reliable information on the number of moments the data possesses (in this case: knows $m\approx 4$), our method adapting to $m$ allows for larger $d$. Finally, Section 6 (cf. also Figure 1 in the Introduction) shows that the trimmed mean of Resende (2024) i) often requires trimming more observations than the sample size $n$, rendering it unimplementable, or ii) its distribution is very poorly approximated by that of $Z$. In contrast, the distribution of $S_{n,W}$ is closely approximated by that of $Z$.
@@
To prove Theorem 2.1, we first show that the order statistics $\hat{\alpha}_{j}$ and $\hat{\beta}_{j}$ can (essentially) be replaced by closely related population quantiles $Q_{\varepsilon_{n},j}$ and $Q_{1-\varepsilon_{n},j}$ of the $X_{1,j}$, such that one can analyze the non-random winsorization functions $\phi_{n,j}(x)\mathrel{\mathop{\ordinarycolon}}=\phi_{Q_{\varepsilon_{n},j},Q_{1-\varepsilon_{n},j}}(x)$ instead of $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(x)$ for all $j=1,\ldots,d$: [^3] A similar control of the order statistics was also used in Lugosi and Mendelson (2021) and Kock and Preinerstorfer (2025). Resende (2024) studied the trimmed mean by relating it to a winsorized mean.
$S_{n,W,j}\approx\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(\tilde{X}_{i,j})-\mu_{j}\bigr]}}_{T_{n,j}}$ $=\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(\tilde{X}_{i,j})-\phi_{n,j}(X_{i,j})\bigr]}}_{I_{n,j,1}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{n,j}(X_{i,j})-\mathbb{E}\phi_{n,j}(X_{i,j})\bigr]}}_{I_{n,j,2}}$
$+\underbrace{\frac{1}{n^{1/2}}\sum_{i=1}^{n}\mathinner{\bigl[\mathbb{E}\phi_{n,j}(X_{i,j})-\mu_{j}\bigr]}}_{I_{n,j,3}}.$
@@
The term $\max_{j=1,\ldots,d}|I_{n,j,1}|$ isolates the effect of the adversarial contamination of the data and $\max_{j=1,\ldots,d}|I_{n,j,3}|$ quantifies how far the winsorized means $\mathbb{E}\phi_{n,j}(X_{1,j})$ are from the population means $\mu_{j}$ of interest. Next, note that $I_{n,j,2}$ is a sample average of bounded i.i.d. random variables. Thus, letting $Z_{n}\sim\mathsf{N}_{d}(0,\Sigma_{\phi_{n}})$ with $\Sigma_{\phi_{n}}$ being the covariance matrix of $I_{n,2}=(I_{n,1,2},\ldots,I_{n,d,2})^{\prime}$, one can apply, e.g., the Gaussian approximation for sums of sub-exponential random vectors from Chernozhukov et al. (2022) to this term to show that
@@ 3 Bootstrap approximations for winsorized means
Similarly to Gaussian approximations for $S_{n}$, cf. the references in the introduction, the one for $S_{n,W}$ in Theorem 2.1 is not directly useful for statistical inference since the covariance matrix $\Sigma$ of the approximating distribution $\mathsf{N}_{d}(0,\Sigma)$ is typically unknown. Because the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 in Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)) shows that for $Z_{1}\sim\mathsf{N}_{d}(0,\Sigma^{(1)})$ and $Z_{2}\sim\mathsf{N}_{d}(0,\Sigma^{(2)})$ with $\min_{j=1,\ldots,d}\Sigma^{(2)}_{j,j}>b$ for some $b>0$, it holds that
$\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(Z_{1}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z_{2}\in H\bigr)}\Bigr\rvert}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\Sigma^{(1)}_{j,k}-\Sigma^{(2)}_{j,k}\bigr\rvert}\Bigr)}^{1/2},$ (6)
@@
In the next section, we construct an estimator $\tilde{\Sigma}_{n}$, say, which is based on suitably winsorized observations $\tilde{X}_{i}$, and for which we establish performance guarantees even for $\bar{\eta}_{n}>0$ and when the $X_{i}$ possess only $m>2$ moments. Our estimator does not depend on unknown population quantities and its performance guarantees “adapt” to the unknown $m$. We stress that Theorem 3.2 below is modular in the sense that it remains valid for any other estimator $\hat{\Sigma}_{n}$ satisfying a bound as in Theorem 3.1.
@@ 3.1 Estimating $\Sigma$
Imposing only $m>2$ moments to exist we now construct an estimator of $\Sigma$ with precision guarantees in the maximal entrywise norm. Set
$\varepsilon_{n}^{\prime}=\lambda_{1}^{\prime}\cdot\overline{\eta}_{n}+\lambda_{2}^{\prime}\cdot\frac{\log(d^{2}n)}{n},\qquad\lambda_{1}^{\prime}\in(1,\infty)\text{ and }\lambda_{2}^{\prime}\in(0,\infty).$ (7)
Like for $\lambda_{1}$ and $\lambda_{2}$, we take $\lambda_{1}^{\prime}$ and $\lambda_{2}^{\prime}$ as fixed throughout.
Writing $\hat{a}_{j}=\tilde{X}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}$$\hat{b}_{j}=\tilde{X}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}$ and $\tilde{\mu}_{n,j}=n^{-1}\sum_{i=1}^{n}\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})$ for $j=1,\ldots,d$, define $\tilde{\Sigma}_{n}$ as the matrix with entries
$\tilde{\Sigma}_{n,j,k}=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\tilde{\mu}_{n,j}\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\tilde{\mu}_{n,k}\bigr]},\qquad 1\leq j,k\leq d.$ (8)
The estimator $\tilde{\Sigma}_{n}$, which is a winsorized analog of the sample covariance matrix, is positive semi-definite and symmetric by virtue of being a Gram matrix. Solely for the purpose of stating the precision guarantee for $\tilde{\Sigma}_{n}$ in Theorem 3.1 below, we introduce the condition
$2\varepsilon_{n}^{\prime}+\frac{\log(d^{2}n)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(d^{2}n)}{n}\Bigr)}^{2}+4\frac{\log(d^{2}n)}{n}\varepsilon_{n}^{\prime}}<1.$ (9)
We stress that (9) is not needed for any of the subsequent Gaussian and bootstrap approximations of this paper to hold, as these will be argued to be trivially true when (9) is violated, cf., e.g., the closing paragraph of the proof of Theorem 3.2. Note also that in the leading special case of $\overline{\eta}_{n}=0$, which is always imposed in the literature when studying Gaussian and bootstrap approximations to the distribution of $S_{n}$, (9) reduces to
$\mathinner{\Bigl(2\lambda_{2}^{\prime}+1+\sqrt{1+4\lambda_{2}^{\prime}}\Bigr)}\frac{\log(d^{2}n)}{n}<1,$
which is often satisfied for $\lambda_{2}^{\prime}$ small.
Remark 3.1.
The condition in (9) is actually just a conservative (simple) sufficient condition for the following milder condition: Writing $A_{+}^{\prime}=1-\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}\in(0,1]$$A_{-}^{\prime}=1+\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}\in[1,\infty)$, and denoting by $W_{0}$ and $W_{-1}$ the principal and lower branch of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)), respectively, (9) can be replaced by
$\varepsilon_{n}^{\prime}\mathinner{\biggl(-A_{+}^{\prime}W_{0}\mathinner{\Bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{+}^{\prime})/A_{+}^{\prime}}\Bigr)}-A_{-}^{\prime}W_{-1}\mathinner{\Bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{-}^{\prime})/A_{-}^{\prime}}\Bigr)}\biggr)}<1.$
By (G.6) of Lemma G.6 in the appendix, the left-hand side of the previous display is upper bounded by the left-hand side of (9), explaining the condition in (9). Note that (9) also implies $\frac{\log(d^{2}n)}{n}<1$, which we repeatedly use in the proofs.
@@
Let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}^{\prime}$ is chosen as in (7) and satisfies (9), then for a constant $C$ depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$, we have
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}\leq\frac{24}{n}.$ (10)
Theorem 3.1 shows that for any $m>2$ it is possible for $\max_{1\leq j,k\leq d}\mathinner{\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\rvert}$ to converge to zero in probability even when $d$ grows exponentially in $n$. Finally, we reiterate that the approximation in Theorem 3.2 below remains valid for any covariance estimator satisfying (10).